OpenStudy (anonymous):

A wooden plate is shaped as a pentagon. The measure of one angle is 40°. The remaining interior angles are obtuse and of equal measure. What is the measure of each obtuse angle in the pentagon? a) 120° b)125° c)110° d)100°

OpenStudy (ganpat):

For a pentagon, the sum of angles = 540.. So given one angle = 40.. and remaining angles are equal.. so 4x + 40 = 540 --->> 4x = 500 x = 125 degree... option B..

OpenStudy (ganpat):

for a polygon, get number of sides.. so number of angles = 180 * (n-2) where, n = number of sides for polygon...

OpenStudy (anonymous):

thanks awesome(:

OpenStudy (ganpat):

W/c..

OpenStudy (anonymous):

?

OpenStudy (ganpat):

Welcome.. :D